# ECGrecover: a Deep Learning Approach for Electrocardiogram Signal Completion

### v1

@@ Abstract.
In this work, we address the challenge of reconstructing the complete 12-lead ECG signal from incomplete parts of it. We focus on two main scenarii: (i) reconstructing missing signal segments within an ECG lead and (ii) recovering missing leads from a single-lead. We propose a model with a U-Net architecture trained on a novel objective function to address the reconstruction problem. This function incorporates both spatial and temporal aspects of the ECG by combining the distance in amplitude between the reconstructed and real signals with the signal trend. Through comprehensive assessments using both a real-life dataset and a publicly accessible one, we demonstrate that the proposed approach consistently outperforms state-of-the-art methods based on generative adversarial networks and a CopyPaste strategy. Our proposed model demonstrates superior performance in standard distortion metrics and preserves critical ECG characteristics, particularly the P, Q, R, S, and T wave coordinates. Two emerging clinical applications emphasize the relevance of our work. The first is the increasing need to digitize paper-stored ECGs for utilization in AI-based applications (automatic annotation and risk-quantification), often limited to digital ECG complete 10s recordings. The second is the widespread use of wearable devices that record ECGs but typically capture only a small subset of the 12 standard leads. In both cases, a non-negligible amount of information is lost or not recorded, which our approach aims to recover to overcome these limitations.
@@ 1. Introduction
An electrocardiogram (ECG) is a diagnostic test broadly used to evaluate cardiac health. It is commonly performed by placing electrodes at specific locations on the patient’s body to capture the heart’s electrical signals. Each electrode records a distinct waveform. The thoracic electrodes (V1-V6) provide specific information about the heart’s electrical signals in areas like the right and left ventricles, while the limb electrodes (I, II, III, aVR, aVL, aVF) offer insights into the heart’s electrical signals in various directions, encompassing left-to-right, superior-to-inferior, and diagonal orientations (Parmet et al., 2003). Healthcare professionals can effectively diagnose and assess cardiac conditions by leveraging the combined information obtained from thoracic and limb electrodes. Indeed, certain diseases may manifest with varied expressions on an ECG depending on the specific lead under examination. For instance, when myocardial infarction occurs in the inferior part of the heart, it produces alterations on the ECG signal that are more visible in the leads II, aVF, and III 111https://ecgwaves.com/topic/localization-localize-myocardial-infarction-ischemia-coronary-artery-occlusion-culprit-stemi/.
@@
While comprehensive diagnosis of potential diseases requires all the leads (in most cases, 12), novel technological advancements have revolutionized how individuals monitor their health. Indeed, ECGs are frequently recorded using wearable devices designed to be worn for longer periods (in contrast with the classical clinical exam, which lasts 10-30 seconds). However, such devices measure only one or several leads in the best cases, e.g., certain smartwatches can record and transmit a single-lead ECG, equivalent to lead I (van der Zande et al., 2023).
@@
Moreover, despite the ongoing shift towards AI and automated ECG analyses, many health establishments, particularly in developing countries, still rely on paper to record and store ECGs (Walker et al., 2009). This poses obvious accessibility issues and makes computer analyses especially challenging. Indeed, paper-stored ECGs are inherently limited when compared with their raw counterparts. In the most common formats, each lead is only partially represented (i.e., only 2.5 or 5 seconds instead of the usual 10-second signal is printed). Although several tools exist (Lence et al., [n. d.]) to digitize the ECG stored on paper, they cannot compensate for the missing signal. Relying on a partial signal for a medical assessment can lead to even more catastrophic consequences: intra-lead variability refers to changes within the same lead of the ECG. Specifically, QT alteration involves fluctuations in the duration of the QT interval. Analyzing a longer segment may provide a clearer picture of QT alteration and other subtle changes that might not be as evident in a shorter segment. For instance, conditions such as bradycardia or extra-systoles require practitioners to analyze more than three heartbeats to determine the presence of the pathology. The significance of QT alterations lies in its association with an increased risk of ventricular arrhythmia, including Torsades-de-Pointes (TdP), which can degenerate into potentially life-threatening ventricular fibrillation (Crotti et al., 2008; Prifti et al., 2021; Salem et al., 2017).
@@
Consequently, reconstructing missing portions of an ECG signal, whether it involves an entire lead or part of it, is a challenging task with implications for both the medical field and computer science.
To approach this problem comprehensively, we formally define it and address it from two perspectives: ECG lead-reconstruction and ECG segment-recovery. In ECG lead reconstruction, the objective is to reconstruct the 11 missing leads using the complete trace of one lead. For ECG segment recovery, the goal is to recover the full signal of 10 seconds for all the leads from only partial segments. Our strategy begins by understanding how the partial signal appears in real-life situations. In Figure 1, an example is provided to illustrate what a 12-lead ECG recording looks like when utilized by clinicians. The names of the leads are indicated in blue, while the actual signals are in black. The recordings from the various leads are concatenated according to the standard 4-channels printing format (e.g., starting with lead I, followed by aVR, V1, and V4). Typically, a 10-second recording of lead II or I is also provided. Lastly, the initial one millisecond of each row serves for amplitude calibration. We utilize a masking approach involving specific masks to replicate real-life scenarii like those illustrated in Figure 1. Given a 12-lead ECG input, the masks are functions that return the same signal with a portion of it preserved and the rest replaced with randomly valued signals. For instance, the 4-channels format mentioned before corresponds to the mask C3 in Figure 2, where the portion of the signal retained is highlighted in green and is referred to as the primer.
Our methodology consists of developing a U-Net-like model, trained with a novel objective function, combining the mean squared error loss (which serves to guide the model to preserve the magnitude of the original ECG signal correctly) with the Pearson loss (which helps the model to maintain the same trend as the original signals). We named the final model ECGrecover. The model was trained on a (4498 ECG) real-life 12-lead, 10-seconds ECGs from a cohort of 990 subjects recruited at
the Pitié-Salpétrière Clinical Investigation Center (Salem et al., 2017). During training, all the masks depicted in Figure 1 are applied to the training subset of the dataset. However, each mask configuration is evaluated separately to quantify its performance during testing.
The main contributions of this work are therefore three-fold:
@@
We identified 17 realistic scenarii where ECG signals may be partially available. Our analysis distinguishes between two primary challenges: reconstructing missing segments within an ECG lead and reconstructing the entire set of 12 leads from one single lead. In addition, a formal definition of the problem is provided (Section 2.1).
@@
(2) We introduce ECGrecover, a model based on the U-Net architecture tailored for segment-recovery and lead-reconstruction tasks. ECGrecover is trained with a novel objective function combining mean squared error and Pearson correlation coefficient. This approach ensures that the reconstructed signals match the original in amplitude and maintain fidelity to their waveform patterns, a crucial aspect for accurate ECG analysis and subsequent clinical applications (Section 2.2).
@@
(3) In close collaboration with clinicians, we extensively evaluated ECGrecover on a dataset composed of real-life ECGs222Code available at https://git.ummisco.fr/open/2024-ecg-recover. (Salem et al., 2017; Prifti et al., 2021). Besides considering classical distortion metrics, this assessment was also focused on the model’s ability to preserve essential features of the ECG signal, including P, Q, R, S, and T waves. Our findings demonstrate ECGrecover’s effectiveness in maintaining the integrity of these signal components (Section 3).
@@ 1.1. Related works
Synthesizing ECG signals involves creating artificial waveforms that resemble real ECG patterns. Generally, ECG synthesis has been studied through the lens of data augmentation (Rahman et al., 2023) aiming to augment the training dataset by generating ‘authentic’-like ECGs conditioned or not on heart disease, thereby to enhancing classifiers performances in arrhythmia predictions (e.g., (Chen et al., 2022; Golany et al., 2020; Golany and Radinsky, 2019)); to improve ECG quality assessment (Zhou et al., 2021), and to improve ECG anomaly detection (Li et al., 2022), among other applications. Existing techniques often involve training deep neural networks such as Variational Autoencoder (VAE) (e.g. (Soleimani and Lobaton, 2022; Kuznetsov et al., 2021)), Generative Adversarial Networks (GAN)(e.g. (Ye et al., 2019; Hazra and Byun, 2020)), or CycleGAN (e.g (Abdelmadjid and Boukadoum, 2022)).
While these authors focus on synthetically generating ECG signals, they do not consider the ECG lead-reconstruction problem or the ECG segment-recovery problem, which can be viewed as a special case of the broader ECG synthesis problem.
@@
ECG lead-reconstruction has been addressed in (Atoui et al., 2010; Wang et al., 2019), where the remaining leads are reconstructed from other leads (I, II, and V2). In (Cho et al., 2020; Lee et al., 2015), the ECGs at limb leads (frontal electrodes) are used to generate the ones at the thoracic leads (precordial electrodes). Nonetheless, the aforesaid studies do not address the scenario where the reconstruction of the remaining leads is required when only a single lead is available, nor when the leads are partially recorded (less than 10 seconds). This is common when patients’ health is monitored using wearable devices that measure only lead I.
@@
An interesting solution is proposed in (Joo et al., 2023), where the authors introduced EKGAN, a conditional GAN based on Pix2Pix (Isola et al., 2017), consisting of two generators and one 1D U-Net (Ronneberger et al., 2015) discriminator, that tackles the reconstruction of 12-lead ECGs from one single-lead (lead I). Prior attempts in synthesizing multiple ECGs from a single-lead ECG were made in (Seo et al., 2022) where the authors proposed a GAN composed of a U-Net generator and a discriminator based on the PatchGAN (Isola et al., 2017) architecture. The model is designed to take 2.5-second ECG segments from lead I as input and reconstruct 2.5-second segments for any of the other leads. This means there are 11 different models, each tailored for reconstructing the ECG segment of a specific lead from the input derived from lead I.
@@
It is worth noting that in both (Joo et al., 2023) and (Seo et al., 2022), the experimental data comprise 10s ECG recordings at 500Hz. However, the evaluation is conducted on different databases: a private one for (Joo et al., 2023) and PTBXL (Wagner et al., 2020) along with a second dataset (Zheng et al., 2020) for (Rautaharju et al., 1981; Seo et al., 2022).
@@
Finally, a naive technique for the ECG segment-recovery problem involves copy-pasting the available signal to fill-up the missing segments, as suggested by Badilini et al. in (Badilini et al., 2005).
@@
2. Methods
In this section, we first formally define the problem of ECG reconstruction. Subsequently, we introduce ECGrecover, our proposed approach, and provide real-life cases of partial ECG we analyze. In Table 1, the list of the symbols used throughout this article.
2.1. Problem formulation
Let be the random variable (r.v.) for which we have realization , where represents the number of leads and the length of the signal in terms of time points. In particular, we will refer to as the set of valid ECGs, meaning they make sense in clinical terms. In this specific context, we say that an ECG is valid if it is also complete meaning that, for all and , , where indicates the specific data point at lead and time point within the ECG matrix while denotes the ECG vector of time points at lead . We call it incomplete otherwise, and we denote the set of incomplete ECGs as .
We aim to develop a model capable of inferring the missing portions of incomplete ECG, whether these are specific segments within leads or entire leads themselves. To achieve this goal, we employ a data-driven approach, training a generative model on masked data to simulate real-life configurations where specific segments (or leads) of the signal are missing: in Figure 2 for each of the twelve leads, the green segments denote the available segments, i.e., the primer, while the red segments denote the part of the ECG that needs to be reconstructed. Specifically, in Figure 2(a), we present configurations with contiguous primers varying in overall signal coverage and lead: primers of 0.8s in C1, primers of 1.6s in C2, primers of 2.5s in C3, primers of 3.3s in C4, and primers of 5.0s in C5. In Figure 2(b), we depict configurations where the full leads must be reconstructed.
The configuration, which is the most considered in the literature, is C, where given the full trace of the first lead, the goal is to reconstruct the remaining ones (Joo et al., 2023). Although in reality, the ECG acquisition machines record the whole signal (12-lead 10-seconds), most of this signal is lost when it is stored in paper ECG format. For instance, some recording machines may capture six leads for 5.0 seconds, as illustrated in C5, or they might record three leads for a shorter span of 2.5 seconds, as in the case of the C3 configuration (Figure 1). It is noteworthy to mention, that our problem definition comprehensively encompasses both scenarii: ECG segment recovery and ECG lead reconstruction, at a higher level of abstraction.
@@ 2.2. Our proposed method
Formally, let be a random realization of i.i.d. valid ECGs according to . Moreover, let be a countable set of indexes corresponding to each of the possible real-life data masks (e.g., as the ones we show in Figure 2). Finally, we let be the set of possible masking functions that, given a valid ECG, outputs an incomplete one. Our training set is defined as
@@
where is the hyperparameter regulating the loss and
@@
The final cost function in Equation 1 accounts for the ECG’s spatial and temporal aspects. Specifically, is the mean squared error loss and ensures an accurate reconstruction of the ECG signal by capturing the average distance in amplitude between the reconstructed and the real signal. On the contrary, is the Pearson loss, which promotes the preservation of specific linear relationships within the signal. A strong correlation (close to 1, in absolute value) signifies that the reconstructed signal varies similarly and simultaneously with the real one. In Section A.1, we detail the performance of the model in the validation set across various values of , particularly focusing on the extreme cases: , i.e., when the loss is exclusively based on , and , i.e., the loss is only.
Similarly to what has been done at training time, at testing time starting from random realization of i.i.d. valid ECGs according to , we apply each of the mask in to create the pair of samples to test. In contrast to the training phase, our evaluation is performed separately for each set of samples generated according to the specific masks under consideration.

### v3

@@ Abstract.
In this work, we address the challenge of reconstructing the complete 12-lead ECG signal from its incomplete parts. We focus on two main scenarios: (i) reconstructing missing signal segments within an ECG lead and (ii) recovering entire leads from signal in another unique lead. Two emerging clinical applications emphasize the relevance of our work. The first is the increasing need to digitize paper-stored ECGs for utilization in AI-based applications, often limited to digital 12 lead 10s ECGs. The second is the widespread use of wearable devices that record ECGs but typically capture only one or a few leads. In both cases, a non-negligible amount of information is lost or not recorded. Our approach aims to recover this missing signal. We propose ECGrecover, a U-Net neural network model trained on a novel composite objective function to address the reconstruction problem. This function incorporates both spatial and temporal features of the ECG by combining the distance in amplitude and sycnhronization through time between the reconstructed and the real digital signals. We used real-life ECG datasets and through comprehensive assessments compared ECGrecover with three state-of-the-art methods based on generative adversarial networks (EKGAN, Pix2Pix) as well as the CopyPaste strategy. The results demonstrated that ECGrecover consistently outperformed state-of-the-art methods in standard distortion metrics as well as in preserving critical ECG characteristics, particularly the P, QRS, and T wave coordinates.
@@ 1. Introduction
An electrocardiogram (ECG) is a diagnostic test broadly used to evaluate cardiac health. It is commonly performed by placing electrodes at specific locations on the patient’s body to capture the heart’s electrical signals. The thoracic electrodes (V1-V6) provide information on the right and left ventricles, while the limb electrodes (I, II, III, aVR, aVL, aVF) offer insights into the heart’s electrical signals in various spatial directions (parmet2003electrocardiograms). Healthcare professionals can effectively diagnose and assess cardiac conditions by leveraging the combined information obtained from these electrodes. However, depending on the specific lead under examination, certain diseases may manifest with varied expressions on an ECG. For instance, when myocardial infarction occurs in the inferior part of the heart, it produces alterations on the ECG signal that are more visible in the leads II, aVF, and III 111https://ecgwaves.com/topic/localization-localize-myocardial-infarction-ischemia-coronary-artery-occlusion-culprit-stemi/.
@@
While a comprehensive disease diagnosis requires all the leads (in most cases, 12), novel technological advancements have revolutionized how individuals monitor their health. ECGs are frequently recorded with wearable devices for longer periods (in contrast with the classical clinical exam, which lasts 10-30 seconds). However, such devices measure only one or a few leads in the best cases, e.g., certain smartwatches can record and transmit a single-lead ECG, equivalent to lead I (van2023using).
@@
Moreover, despite the ongoing shift towards AI and automated ECG analyses, the majority of healthcare establishments around the world, particularly in developing countries, still rely on paper to record and store ECG (walker2009low). This poses obvious accessibility issues and makes computer-based analyses especially challenging. In the most common formats, each lead is only partially represented (i.e., only 2.5 or 5 seconds). Although several tools exist (lence_automatic_2023) to digitize the ECG stored on paper, they cannot compensate for the missing signal. Relying on a partial signal for a medical assessment can lead to even more catastrophic consequences. For example, QT alteration involves fluctuations in the duration of the QT interval. Analyzing a longer segment may provide a clearer picture of QT alteration and other subtle changes that might not be visible in shorter segments. Conditions such as bradycardia or extrasystoles require practitioners to analyze more than three heartbeats to determine the presence of the pathology. The significance of QT alterations lies in its association with an increased risk of ventricular arrhythmias, including Torsades-de-Pointes (TdP), which can degenerate into potentially life-threatening ventricular fibrillation (crotti2008congenital; prifti2021deep; salem2017genome)
@@
Consequently, reconstructing missing portions of an ECG signal, whether it involves an entire lead or part of it, is a challenging task with implications for both the medical field and computer science. To approach this problem, we formally define it and address it from two perspectives: ECG lead-reconstruction and ECG segment recovery. In ECG lead reconstruction, the purpose is to reconstruct the 11 missing leads using the full trace of only one lead, while for ECG segment recovery, the goal is to recover the entire 10-second signal for all the leads from only partial existing segments.
In Figure 1, we show a 12-lead ECG recorded in paper format. The recordings from the various leads are concatenated according to the standard 3x4 printing format (3 partial channels with 4 partial leads each, starting with lead I, followed by leads aVR, V1, and V4, and a fourth channel containing a complete lead, either lead II or I). Lastly, the reference pulses are used for amplitude calibration at the beginning of each row. We proposed a masking approach involving specific masks to replicate real-life scenarios like those illustrated in Figure 1. Given a 12-lead ECG input, the masks are functions that return the same signal, with parts of it replaced by random values drawn from a uniform distribution. For instance, the 3x4 format corresponds to the C3 mask in Figure 2, where the portion of the signal retained is in green, which we refer to as the primer.
Our approach involved developing a U-Net-like neural network model, trained with a novel original objective function, combining the mean squared error loss with the Pearson correlation loss. The former serves to guide the model in preserving the magnitude of the original ECG signal, while the latter allows the model to maintain synchronization with the original signal. We named this approach ECGrecover. The final model was trained on a dataset comprising 4498 real-life 12-lead, 10-second ECGs from a cohort of 990 healthy volunteers (salem2017genome) who were administered 80mg of Sotalol, a QT prolonging drug. During training, all the masks depicted in Figure 1 were applied to each ECG of the dataset. However, each mask was applied separately to the testing set during the evaluation.
The main contributions of this work are threefold:
@@
We identified 17 realistic scenarios where ECG signals may be partially available. Our analysis differentiates between two cases: reconstructing missing segments within an ECG lead and reconstructing the entire set of 12 leads from one single lead. In addition, a formal definition of the problem is provided (Section 2.1).
@@
(2) We introduce ECGrecover222Code available at https://github.com/UMMISCO/ecgrecover, a U-Net model for segment-recovery and lead-reconstruction tasks. ECGrecover is trained with a novel objective function combining the mean squared error and the Pearson correlation to ensure the reconstructed signal matches the original in amplitude and preserves synchronized waveform patterns, crucial for clinical applications (Section 2.2).
@@
(3) In close collaboration with expert cardiologists, we extensively evaluated ECGrecover on a dataset composed of real-life ECG (salem2017genome; prifti2021deep). Besides considering classical distortion metrics, this assessment also focused on the model’s ability to preserve clinically relevant signal components of the ECG signal, such as QT distance (Section 3).
@@ 1.1. Related works
Synthesizing ECG signals involves creating artificial waveforms that resemble real ECG patterns. Generally, ECG synthesis has been studied through the lens of data augmentation (rahman2023systematic) aiming to increase the size of the training dataset by generating ”authentic”-like ECGs conditioned or not on heart pathologies, to enhance classifier performance in arrhythmia predictions (e.g., (chen2022me; golany2020improving; golany2019pgans)); to improve ECG quality assessment (zhou2021electrocardiogram), and to improve ECG anomaly detection (li2022contrastive), among other applications. Existing techniques often involve training deep neural networks such as Variational Autoencoders (VAE) (e.g. (soleimani2022enhancing; kuznetsov2021interpretable)), Generative Adversarial Networks (GAN) (e.g. (ye2019ecg; hazra2020synsiggan)), or CycleGAN (e.g (abdelmadjid2022neural)).
@@
ECG completion has been addressed in (atoui2010novel; wang2019novel), where the remaining leads are reconstructed from three other complete leads (I, II, and V2). In (cho2020artificial; lee2015reconstruction), the limb leads (frontal electrodes) are used to generate the thoracic leads (precordial electrodes). Nonetheless, these studies do not address the scenario where the reconstruction of the remaining leads is based only on a single lead, nor when the leads are partially recorded (less than 10 seconds).
@@
Another solution is proposed in (joo2023twelve), where the authors introduced EKGAN, a conditional GAN based on Pix2Pix (isola2017image), consisting of two generators and one 1D U-Net discriminator (ronneberger2015u), that tackles the reconstruction of 12-lead ECGs from one single lead (lead I). Prior attempts in synthesizing multiple ECGs from a single-lead ECG were made in (seo2022multiple) where the authors proposed a GAN composed of a U-Net generator and a discriminator based on the PatchGAN (isola2017image) architecture. The model is designed to take 2.5-second ECG segments from lead I as input and reconstruct 2.5-second segments for any other leads. The authors trained 11 different models, each tailored for reconstructing the ECG signal of a specific lead from the input derived from lead I.
@@
It is worth noting that in both (joo2023twelve) and (seo2022multiple), the experimental data comprise 10s ECG recordings at 500Hz. However, the evaluation is conducted on different databases: a private one for (joo2023twelve) and PTBXL (wagner2020ptb) along with a second dataset (zheng202012) for (rautaharju1981exploitation; seo2022multiple).
@@
Finally, a more naive technique for the ECG segment-recovery problem involves copy-pasting the available signal to fill up the missing segments, as suggested by Badilini et al. in (badilini2005ecgscan).
Considering a broader context beyond just ECGs, Alcaraz et al. (alcaraz2022diffusion) propose a time series imputation model that integrates the principles of classical conditional time diffusion models with the methodologies of structured state space models. Unlike our setting, which focuses on 17 realistic cases, they focus on scenarios not necessarily connected to our analysis. Specifically, they address random missing, where zero entries of an imputation mask are sampled randomly across all channels; random block missing (RBM), where one segment per channel is sampled as the imputation target; blackout missing (BM), where a single segment is missing across all channels; and time series forecasting (TF), a special case of BM where the imputation region spans time steps at the end of the sequence. While Alcaraz et al. provide experiments on the PTB-XL datasets, their method cannot be fairly compared to ours due to key differences. Their ECGs are downsampled from 500 Hz to 100 Hz, which could shift data distributions significantly. Additionally, their model’s training time is considerably longer at 36 hours, affecting its practicality and scalability. Finally, we would like to point out that, in our configurations, the models must reconstruct up to 92% of the original signal in the worst-case scenario (C1), whereas the proposed schemes in (alcaraz2022diffusion) only require up to 50% of the signal to be generated each time (RBM setting).
2. Methods
In this section, we formally define the problem of ECG reconstruction. Subsequently, we introduce ECGrecover, our proposed approach, and extensively evaluate it while comparing it with state-of-the-art methods using relevant real-life ECG datasets. In Table 1, the list of the symbols used throughout this article.
2.1. Problem formulation
Let be the random variable (r.v.) for which we have realization , where represents the number of leads and the length of the signal in terms of sample points. We will refer to as the set of valid ECGs, meaning they make sense in clinical terms. In this specific context, we say that an ECG is valid if it is also complete i.e., for all and , , where indicates the specific data point at lead and time point within the ECG matrix, while denotes the ECG vector of time points at lead . We call it incomplete otherwise, and we denote the set of incomplete ECGs as .
@@ 2.1. Problem formulation
We aimed to develop a model capable of inferring the missing portions of incomplete ECG, whether these are specific segments within leads or the entire leads themselves. To achieve this goal, we employed a data-driven approach, training a generative model on masked data to simulate real-life configurations where specific segments (or leads) of the signal are missing: in Figure 2 for each of the twelve leads, the green segments denote the available segments, i.e., the primer, while the red segments denote the part of the ECG to be reconstructed. Specifically, in Figure 2(a), we present configurations with contiguous primers varying in overall signal coverage and lead: primers of 0.8s in C1, primers of 1.6s in C2, primers of 2.5s in C3, primers of 3.3s in C4, and primers of 5.0s in C5. In Figure 2(b), we depict configurations where the full leads must be reconstructed. The configuration the most considered in the literature is C, where given the full trace of the first lead, the goal is to reconstruct the remaining ones (joo2023twelve). Although, in reality, the ECG acquisition machines record the whole signal (12 lead 10 seconds), most of this signal is lost when stored in paper ECG format or as a digital PDF file. For instance, some recording machines may capture six leads for 5 seconds, as illustrated in C5, or they might record three leads for a shorter span of 2.5 seconds, as in the case of the C3 configuration (Figure 1). Finally, we highlight that at a higher level of abstraction, our problem definition comprehensively encompasses both ECG segment recovery and ECG lead reconstruction.
@@ 2.2. Our proposed method
Formally, let be a random realization of i.i.d. valid ECG according to . Moreover, let be a countable set of indexes corresponding to each of the possible real-life data masks (e.g., as the ones we show in Figure 2). Finally, we let be the set of possible masking functions that, given a valid ECG, outputs an incomplete one. Our training set is defined as
@@
where is the hyper-parameter regulating the loss and
@@
The final cost function in Equation 1 accounts for the ECG’s spatial and temporal aspects. Specifically, is the mean squared error loss and ensures an accurate reconstruction of the ECG signal by capturing the average distance in amplitude between the reconstructed and the real signal. On the contrary, is the Pearson loss, which promotes the preservation of specific linear relationships within the signal. Specifically, as the imputation improves, as (strong correlation). In Section A.1, we detail the performance of the model in the validation set across various values of , particularly focusing on the extreme cases: , i.e., when the loss is exclusively based on , and , i.e., the loss is only. Moreover, we illustrate the loss evolution during training when in Section A.2.
Similarly to training, at testing starting from random realization of i.i.d. valid ECG according to , we apply each one of the masks in to create the pair of samples to test. In contrast to the training phase, our evaluation is performed separately for each set of samples generated according to the specific masks under consideration to evaluate generalization across configurations.